# myboxenhanced,colback=blue!5!white,colframe=blue!75!black,boxrule=0.6mm,fontupper=

### v1

@@ 5.1 Abstraction Pattern
Actions Latent
Factored
Relational
Modular
Rewards Latent
Factored
Dynamics Latent
Factored
Modular
Generalization.
State abstractions are a general choice for improving generalization performance by capturing shared dynamics across MDPs into abstract state spaces using methods such as Invariant Causal Prediction (?, ?), Bisimulation (?, ?), Free Energy Minimization (?), and disentanglement (?, ?).
Value functions can serve as temporal abstractions for shared dynamics in Multi-task Settings. Successor Features (SF) (?, ?) exploit latent reward and dynamic decompositions by using value functions as an abstraction. Subsequent works have combined them with Generalized Policy Iteration (?) and Universal Value Function Approximators (?, ?). Factorization in value functions, other the other hand, can help improve sample efficiency and generalization both (?, ?).
Relational abstractions contribute to generalization by incorporating symbolic spaces into the RL pipeline. These can particularly help with incorporating planning approaches in hierarchical frameworks (?, ?). Additionally, relational abstractions can help abstract away general aspects of a collection of MDPs, thus allowing methods to learn generalizable Q-values over abstract states and actions that can be transferred to new tasks (?) or develop methods specifically for graph-structured spaces (?, ?).
Abstractions can additionally enable generalization in hierarchical settings by compressing state spaces (?), abstract automata (?, ?), or templates of dynamics across tasks (?).
Sample Efficiency
Latent variable models improve sample efficiency across the RL pipeline. Latent state abstractions can improve sample efficiency in Model-based RL (?) and also help improve the tractability of policy learning over options in HRL (?). In model-free tasks, these can also be learned as inverse models for visual features (?) or for control in a latent space (?). Latent transition models demonstrate efficiency gains by capturing task-relevant information in noisy settings (?), by preserving bisimulation distances between original states (?), or by utilizing factorized abstractions (?).
Latent action models can expedite convergence of policy gradient methods such as REINFORCE (?)) by shortening the learning horizon in stochastic scenarios like dialog generation ? (?). Action embeddings, on the other hand, can help reduce the dimensionality of large action spaces ? (?)
Safety and Interpretability
Relational abstractions are a very good choice for interpretability since they capture interactionally complex decompositions. The combination of object-centric representations and learned abstractions can help add transparency (?) while symbolic interjections, such as tracking the relational distance between objects, can help improve performance  (?).
State and rewards abstractions can help with safety. Latent states can help to learn safe causal inference models by embedding confounders (?) On the other hand, meshes (?, ?) help benchmark metrics such as robustness in a learned policy.
5.2 Augmentation Pattern
The augmentation pattern treats $X$ and $z$ as separate input entities for the action-selection mechanism. The combination can range from the simple concatenation of structural information to the state or actions to more involved methods of conditioning policy or value functions on additional information. Crucially, the structural information neither directly influences the optimization procedure nor changes the nature of $X$. It simply augments the already existing entities. In this view, abstractions that are learned in an auxiliary manner and concatenated to states, actions, or models can also be considered augmentations since the original entity remains unchanged.
For the taxi example, one way to apply the augmentation pattern would be by conditioning the policy on additional information, such as the time of day or day of the week. This information could be useful because traffic conditions and passenger demands can vary depending on these factors. However, augmentations can increase the complexity of the policy, and care needs to be taken to ensure that the policy does not overfit the additional information. Due to this, this pattern is generally not explored to its fullest extent. While we see usages of augmentations equitably for most use cases in Figure 5, the number of papers utilizing this pattern still falls short in comparison to more established techniques, such as abstraction. In the next paragraphs, we delineate three kinds of augmentations in the surveyed work.
Space Type Efficiency Generalization Interpretabiltiy Safety
Goals Latent
Factored
Relational
Modular
States Latent
Factored
Relational
Modular
Actions Latent
Relational
Modular
Rewards Factored
Dynamics Latent
Factored
Policies Modular
Context-based Augmentations.
Contextual representations of dynamics (?, ?) and goal-related information (?, ?) can help with generalization and sample efficiency by exposing the agent to the necessary information for optimality. Goal augmentations additionally allow interpretable mechanisms for specifying goals (?). On the other hand, augmentation of meta-learned latent spaces to the normal state can promote temporally coherent exploration across tasks (?). Action histories (?) can directly help with sample efficiency, and action relations (?, ?) contribute to generalization over large action sets.
Language Augmentations.
Language can explicitly capture relational metadata in the world. Latent language interpretation models (?) can utilize the compositionality of language to achieve better exploration and generalization to different relational settings, as represented by their language descriptions. On the other hand, goal descriptions (?) can help hierarchical settings by exploiting semantic relationships between different subtasks and producing better goals for lower-level policies. Augmentations can additionally help make existing methods more interpretable through methods such as ? (?) by guiding search over approximate policies written in human-readable formats.
Control Augmentations.
Augmentations can additionally help with primitive control, such as multi-level control in hierarchical settings. Augmenting internal latent variables conditioned on primitive skills (?, ?, ?) can help tackle sample efficiency in hierarchical settings. Augmentations can also help morphological control (?) through methods such as ? (?) that model the different limbs as individual agents that need to learn to join together into a morphology to solve a task.
5.3 Auxiliary Optimization Pattern
This pattern uses structural side information to modify the optimization procedure. This includes methods involving contrastive losses, reward shaping, concurrent optimization, masking strategies, regularization, baselining, etc. However, given that the changes in the optimization can go hand-in-hand with modifications of other components, this pattern shares methods with a lot of other patterns. For example, contrastive losses can be used to learn state abstractions. Similarly, a learned model can be utilized for reward shaping as well. Thus, methods that fall into this category simultaneously utilize both patterns.
In the case of the taxi, reward shaping could help the policy to be reused for slight perturbances in the city grid, where the shaped reward encourages the taxi to stay near areas where passengers are frequently found when it does not have a passenger. It is crucial to ensure that the modified optimization process remains aligned with the original objective, i.e., there needs to exist some form of regularization that controls how the modification of the optimization procedure respects the original objective. For reward shaping techniques, this amounts to the invariance of the optimal policy under the shaped reward (?). For auxiliary objectives, this manifests in some form of entropy (?) or divergence regularization (?). Constraints ensure this through recursion (?), while baselines control the variance of updates (?). The strongest use of constraints is in the safety literature, where constraints either help control the updates using some safety criterion or constrain the exploration. Consequently, in Figure 5, we see that the auxiliary optimization pattern peaks in its proclivity towards addressing safety. In the following paragraphs, we cover methods that optimize individual aspects of the optimization procedure, namely, rewards, learning objectives, constraints, and parallel optimization.
Space Type Efficiency Generalization Interpretabiltiy Safety
Goals Latent
Relational
Factored
Modular
States Latent
Factored
Relational
Modular
Actions Latent
Factored
Modular
Rewards Factored
Dynamics Latent
Factored
Relational
Policy Space Latent
Reward Modification.
Reward shaping is a very common way to incorporate additional information into the optimization procedure. Methods can gain sample efficiency by exploiting modular and relational decompositions through task descriptions (?), or goal information from a higher level policy with off-policy modification to the lower level transitions (?). Histories of rewards (?) can help learn symmetric relationships between states and, thus, improve the selection procedure for states in a mini-batch for optimization. Factorization of states and rewards into endogenous and exogenous factors (?), on the other hand, helps with safety and sample efficiency through reward corrections.
Extrinsic Rewards can also be used to guide the exploration process. Symbolic planning with relational representations can be used to interact with a primitive learning policy through extrinsic rewards in hierarchical settings, thus, adding interpretability while directly impacting the exploration through the extrinsic reward (?). Alternatively, additional reward sources can help determine the quality of counterfactual trajectories, which can help explain why an agent took certain kinds of actions (?). Additionally, running averages can rewards be used to adaptively tune exploration parameters for heterogeneous action spaces (?)
On the other hand, intrinsic rewards can specifically help with exploration in sparse reward environments. Latent decompositions can help improve such methods by directly impacting the exploration. Language abstractions can serve as latent decompositions that can be separately used for exploration (?). Alternatively, geometric structures can provide a way to compare state embeddings and provide episodic bonuses (?).
Auxiliary Learning Objectives.
Skill-based methods transfer skills between agents with different morphology by learning invariant subspaces and using those to create a transfer auxiliary objective (through a reward signal) (?), or an entropy-based term for policy regularization (?). In hierarchical settings, discovering appropriate sub-tasks (?) can be a highly sample-inefficient process. ? (?) tackle this by composing values of the sub-trajectories under the current policy, which they subsequently use for behavior cloning. Latent decompositions can help with robustness and safety when used for some form of policy regularization (?). Auxiliary losses, which usually help with generalization, can also be a very good entry point for human-like inductive biases (?). Metrics inspired by the geometry of latent decompositions can help with learning optimal values in multi-task settings (?).
Constraints and Baselines.
Constrained optimization is commonplace in Safe RL, and incorporating structure can help improve the sample efficiency of such methods while making them more interpretable. Factorization of states into safe and unsafe states can help develop persistent safety conditions (?), or language abstractions (?). Recursive constraints (?) can help explicitly condition the optimization on a latent subset of safe actions using factored states. Restricting the exploration of options to non-risky states can help incorporate safety in hierarchical settings as well (?). Factorized actions can also help improve the sample efficiency of policy gradient methods through baselining (?, ?), offline methods through direct value conditioning (?), and value-based planning through matrix estimation (?)
Methods can also directly incorporate expert domain knowledge directly in the action selection mechanism for safety and interpretability (?), or for directed exploration to improve sample efficiency (?). Hierarchical settings can benefit from latent state decompositions incorporated via modification of the termination condition (?). Additionally, state-action equivalences can help scale up Q-learning to large spaces through factorization (?).
Concurrent Optimization.
Parallelizing optimization using structural decompositions can specifically help with sample efficiency. Factored MDPs are a very good way to model factors that influence the content presented to users and can be used for ensembling methods in a parallel regime (?). In a similar vein, factored rewards in hierarchical settings can help decompose Multi-task problems into a linear combination of individual task MDPs (?). Alternatively, discretizing continuous sub-actions in multi-dimensional action spaces can help extend the MDP for each sub-action to an undiscounted lower-level MDP, modifying the backup for the Q values using decompositions (?). Relational decompositions can additionally help with masking strategies for Factored Neural Networks (?).
5.4 Auxiliary Model Pattern
This pattern represents using the structural information in a model. In using the term model, we specifically refer to methods that utilize a model of the world to generate experiences, either fully or partially. This notion allows us to capture a range of methods, from ones using full-scale world models to generate rewards and next-state transitions to ones that use these methods to generate full experience sequences. In our categorization, we specifically look at how the structure is incorporated into such models to help generate some parts of learning experiences.
Our taxi agent could learn a latent model of city traffic based on past experiences. This model could be used to plan routes that avoid traffic and hence reach destinations faster. Alternatively, the agent could learn an ensembling technique to combine multiple models, each of which model-specific components of the traffic dynamics. With models, there is usually a trade-off between model complexity and accuracy, and it is essential to manage this carefully to avoid overfitting and maintain robustness. To this end, incorporating structure helps in making the model-learning phase more efficient while allowing reuse for generalization. Hence, in Figure 5, we see that the auxiliary model pattern shows a strong proclivity to utilizing structure for sample efficiency. In the following paragraphs, we explicitly discuss models that utilize decompositions and models used for creating decompositions.
Space Type Efficiency Generalization Interpretabiltiy Safety
Goals Factored
Relational
Modular
States Latent
Factored
Relational
Modular
Actions Latent
Factored
Relational
Modular
Rewards Latent
Factored
Dynamics Latent
Factored
Relational
Modular
Models with structured representations.
? (?) utilize factored decomposition for state space to demonstrate the benefits of model-based methods in combinatorially complex environments. In a similar vein, the dreamer models (?, ?) utilize latent representations of pixel-based environments.
Object-oriented representation for states can help bypass the need to learn latent factors using CNNs in MBRL (?) or as random variables whose posterior can be refined using NNs (?). Graph (Convolutional) Networks (?) can capture rich higher-order interaction data, such as crowd navigation ? (?), or invariances (?). Action equivalences can help learn latent models (Abstract MDPs) (?) for planning and value iteration.
Models for task-specific decompositions.
Another way to utilize decompositions in models is to capture task-specific decompositions. Models that capture some form of relevance, such as observational and interventional data in Causal RL (?), or task-relevant vs. irrelevant data (?) can help with generalization and sample efficiency gains. Latent representations help models capture control-relevant information (?) or subtask dependencies (?).
Models for safety usually incorporate some measure of cost to abstract safe states (?), or unawareness to factor states and actions (?). Alternatively, models can also directly guide exploration mechanisms through latent causal decompositions (?) and state subspaces (?) to gain sample efficiency. Generative methods such as CycleGAN (?) are also very good ways to use Latent models of different components of an MDP to generate counterfactual trajectories (?)
5.5 Warehouse Pattern
This pattern uses structural information to create a database of entities that can then be combined to achieve a specific objective. These can be learned policies and value functions or even models. Given the online nature of such methods, they are often targeted toward continual and life-long learning problems. The inherent modularity in such methods often leads them to focus on knowledge reuse as a central theme.
The taxi from our running example could maintain a database of value functions or policies for different parts of the city or at different times of the day. These could be reused as the taxi navigates through the city, making learning more efficient. While warehousing generally can improve efficiency, it has primarily been explored through the skills and options framework for targeting generalization. An important consideration in warehousing is to manage the size and diversity of the warehouse so as to avoid biasing the learning process too much toward past experiences.
So far, the warehousing pattern seems to be applied to sample efficiency and generalization. However, warehousing also overlaps with interpretability since the stored data can be easily used to analyze the agent’s behavior and understand the policy for novel scenarios. Consequently, these objectives are equitably distributed in Figure 5.
Space Type Efficiency Generalization Interpretabiltiy Safety
Goals Factored
Relational
Modular
States Latent
Factored
Modular
Actions Latent
Modular
Rewards Factored
Dynamics Latent
Factored
Modular
Policies Latent
Modular
Policy Warehousing.
Policy subspaces (?) is a relatively new concept that utilizes shared latent parameters in policies to learn a subspace that can be subsequently combined linearly to create new policies. Extending these subspaces by warehousing additional policies naturally extends them to continual settings (?).
Using goals and rewards, task factorization endows warehousing policies and Q values in multi-task lifelong settings. The multi-task lifelong problem can be treated as a relationship graph between existing tasks generated from latent space (?). On the other hand, methods such as ? (?) factor MDPs into agent-specific and task-specific degrees of variation, for which individual modules can be trained. Disentanglement using variational encoder-decoder models (?) can help control morphologically different agents by factorizing dynamics into shared and agent-specific factors. Additionally, methods such as ? (?) partition the agent’s problem into interconnected sub-agents that learn local control policies.
Methods that utilize the skills framework effectively warehouse learned primitives, similar to how options warehouse associated policies in HRL. These can subsequently be used for maximizing mutual information in lower layers (?), sketching together a policy (?), diversity-seeking priors in continual settings (?), or for partitioned states spaces (?). In a similar vein, ? (?) apply the warehouse pattern on a latent embedding space, learned using auxiliary optimization.
Decomposed Models.
Decompositions that inherently exist in models lead to approaches that often ensemble multiple models that individually reflect different aspects of the problem. Ensemble methods such as ? (?) capture the dynamics in individual modules that sparsely interact and use attention mechanisms (?). Ensembling dynamics can also help with few-shot adaptation to unseen MDPs (?). Factored models can also be combined with relational decompositions to bind actions to object-centric representations ? (?). Latent representations in hierarchical settings (?) can additionally improve the sample inefficiency of Deep Option Critic (?).
5.6 Environment Generation Pattern
This pattern uses structural information to create task, goal, or dynamics distributions from which MDPs can be sampled. This subsumes the idea of procedurally generated environments while additionally incorporating methods that use auxiliary models inducing structure in the environment generation process. The decomposition is reflected in the aspects of the environment generation that are impacted by the generative process, such as dynamics, reward structure, state space, etc. Given the online nature of this pattern, methods in this pattern end up addressing curriculum learning in one way or another.
In the taxi example, a curriculum of tasks could be generated, starting with simple tasks (like navigating an empty grid) and gradually introducing complexity (like adding traffic and passengers with different destinations). Ensuring that the generated MDPs provide good coverage of the problem space is crucial to avoid overfitting to a specific subset of tasks. This necessitates additional diversity constraints that need to be incorporated into the environment generation process. Structure, crucially, provides additional interpretability and controllability in the environment generation process, thus, making benchmarking easier than methods that use unsupervised techniques (?).
Space Type Efficiency Generalization Interpretabiltiy Safety
Goals Relational
Modular
States Latent
Factored
Relational
Rewards Latent
Factored
Dynamics Latent
Factored
Relational
Modular
The compositional nature of learning problems can be modeled using rule-based grammar. ? (?) particularly utilize this to impact the transition dynamics and generate environments. This allows them to train agents with an implicit compositional curriculum. This is further used by ? (?) in their auxiliary optimization procedure. Another way to capture task dependencies is through latent graphical models, which can be used to generate the state-space, reward functions, and transition dynamics (?, ?).
Latent dynamics models allow simulating task distributions, which can help with generalization (?). Clustering methods such as (?), on the other hand, explore task similarities by meta-learning a clustering method through an exploration policy. In a way, they recover a factored decomposition on the task space where individual clusters can be further used for policy adaptation.
5.7 Explicitly Designed
This pattern encompasses all methods where the inductive biases manifest in specific architectures or setups that reflect the decomposability of the problem that they aim to utilize. Naturally, this includes highly specific Neural architectures, but it also easily extends to other methods like sequential architectures to capture hierarchies, relations, etc. Crucially, the usage of structural information is limited to the specificity of the architecture and not any other part of the pipeline. In the case of the taxi, a neural architecture could be designed to process the city grid as an image and output a policy. Techniques like convolutional layers could be used to capture the spatial structure of the city grid. Different parts of the network could be specialized for different subtasks, like identifying passenger locations and planning routes. However, this pattern involves a considerable amount of manual tuning and experimentation, and it’s critical to ensure that these designs generalize well across different tasks. Designing specific neural architectures can provide better interpretability, enabling the ability to decompose different components and simulate them independently. Consequently, this pattern shows the highest proclivity to interpretability, with Generalization being a close second in Figure 5.
Space Type Efficiency Generalization Interpretabiltiy Safety
Goals Factored
Relational
States Latent
Factored
Relational
Modular
Actions Latent
Factored
Relational
Rewards Latent
Factored
Dynamics Latent
Factored
Relational
Policies Relational
Modular
@@ 5.7 Explicitly Designed
Specialized architectures can also help capture invariance in the problem. Symbolic Networks (?, ?) train a set of shared parameters for Relational MDPs by first converting them to Graphs and then capturing node embeddings using Neural Networks. Homomorphic Networks (?) capture symmetry into specialized MLP and CNN architectures. An alternate approach to incorporating symmetry is through basis functions (?).
Attention mechanisms can explicitly capture entity-factored scenarios (?, ?, ?). Relational and Graph Networks can capture additional relational inductive biases explicitly. Linear Relation Networks (?) provides such an architecture that scales linearly with the number of objects. Graph networks have also been used to model an agent’s morphology in embodied control explicitly (?, ?).
@@
A class of methods combines the best of both worlds by capturing invariance in additional specialized modules. Such modules can capture relational structure in semantic meaning (?), relational encoders for auxiliary models (?), or specialized architectures for incorporating domain knowledge (?).
@@ 6 Open Problems in Structured Reinforcement Learning
Having discussed our patterns-oriented framework for understanding how to incorporate structure into the RL pipeline, we now turn to connect our framework with existing sub-fields of RL. We examine existing paradigms in these sub-fields from two major perspectives: Scalability and Robustness. These dimensions serve as a canvas upon which we can position and understand different RL paradigms in sub-fields such as Offline RL, Unsupervised RL, Foundation Models in RL, Partial observability, Big Worlds, Automated RL, and Meta-RL.
Sparse data scenarios require more intelligent ways to make use of limited experiences, while abundant data scenarios might suffer from the quality of data since they might be generated from noisy and often unreliable sources.
@@
measures how well methods scale with the increasing problem complexity in terms of the size of the state and action spaces, complex dynamics, noisy reward signals, and longer task horizons. On one hand, methods might specifically require low dimensional spaces and might not scale so well with increasing the size of these spaces, and on the other, some methods might be overkill for simple problems but better suited for large spaces.
@@
In the following subsections, we cover sub-fields of RL that lie at different points of the 3D space of Data Dependency, Scalability, and Robustness. We introduce the existing paradigms and the current challenges for each of these sub-fields. We then present some example manners in which our framework can bolster further research and practice in these fields. Finally, we collate this discussion into takeaways that can be combined into specific design patterns utilizing our framework.
@@ 6.1 Offline RL
Offline Reinforcement Learning (also known as batch RL) (?) involves learning from a fixed dataset without further interaction with the environment. This approach can benefit in scenarios where active exploration is costly, risky, or infeasible. Consequently, such methods are highly data-dependent due to their reliance on the collected dataset, and they do not generalize well due to the limitations of the pre-collected data. The three dominant paradigms in Offline RL – Behavior Cloning (?), Q-Learning (?), and Sequence Modelling (?) – uniformly degrade in performance as the state-space increases (?). Offline RL also comes with its own challenges, including overcoming distributional shifts and exploiting the available dataset effectively. Structural decomposition can play a crucial role in addressing these challenges in the following ways:
@@
Task decomposition allows learning individual policies or value functions for different subtasks, which could potentially leverage the available data more effectively. For example, modular decomposition through warehousing separate policies for individual modules using the corresponding subset of the data might be more sample-efficient than learning a single policy for the entire task. Task decompositions, thus, open up new avenues for developing specialized algorithms that effectively learn from limited data about each subtask while balancing the effects of learning different subtasks. Practitioners can leverage such decompositions to maximize the utility of their available datasets by training models that effectively handle specific subtasks, potentially improving the overall system’s performance with the same dataset.
@@
The structural information could potentially help mitigate the effect of distributional shifts. For instance, if some factors are less prone to distributional shifts in a factored decomposition, we could focus more on those factors during learning. This opens up venues for gaining theoretical insights into the complex interplay of structural decompositions, task distributions, and policy performance. On the other hand, practical methods for environments where distributional shifts are common could leverage structural decomposition to create more robust RL systems.
@@
Structural decomposition can be used to define auxiliary tasks that facilitate learning from the dataset. For instance, in a relational decomposition, we could define auxiliary tasks that involve predicting the relationships between different entities, which could help in learning a useful representation of the data. Using the proposed framework, researchers can explore how to define meaningful auxiliary tasks that help the agent learn a better representation of the environment. This could lead to new methods that efficiently exploit the available data by learning about these auxiliary tasks. Practitioners can design auxiliary tasks based on the specific decompositions of their problem. For example, if the task has a clear relational structure, auxiliary tasks that involve predicting the relations between different entities can potentially improve the agent’s understanding of the environment and its overall performance.
{mybox}
Offline RL Patterns
  • Use a Factored or Relational decomposition to create abstractions that can help with distribution shift and auxiliary interpretability.
  • Implement a Modular design with each module targeting a specific sub-problem, which would improve Scalability.
  • Employ policy reuse by warehousing policies learned for sub-problems across tasks.
  • If sufficient interaction data is available, employ data augmentation strategies for counterfactual scenarios using latent models.

### v3

@@ 6.7 Explicitly Designed
Actions Latent ? (?)
Factored ? (?) ? (?)
Relational ? (?) ? (?)
Rewards Latent ? (?)
Factored ? (?)
Dynamics Latent ? (?) ? (?), ? (?)
Factored ? (?),? (?)
Relational ? (?)
Policies Relational ? (?), ? (?) ? (?) ? (?)
Modular ? (?) ? (?),? (?)
Survey of literature utilizing the explicitly designed pattern
@@
Specialized architectures also help capture invariance in the problem. Symbolic Networks (?, ?) train a set of shared parameters for Relational MDPs by first converting them to Graphs and then capturing node embeddings using Neural Networks. Homomorphic Networks (?) capture symmetry into specialized MLP and CNN architectures. An alternate approach to incorporating symmetry is through basis functions (?).
Attention mechanisms explicitly help capture entity-factored scenarios (?, ?, ?). Relational and Graph Networks capture additional relational inductive biases explicitly. Linear Relation Networks (?) provides an architecture that scales linearly with the number of objects. Graph networks have also been used to model an agent’s morphology in embodied control explicitly (?, ?).
@@
These are a class of methods that combines the best of both worlds by capturing invariance in additional specialized modules. Such modules capture relational structure in semantic meaning (?), relational encoders for auxiliary models (?), or specialized architectures for incorporating domain knowledge (?).
@@ 7 Open Problems in Structured Reinforcement Learning
Having discussed our patterns-oriented framework for understanding how to incorporate structure into the RL pipeline, we now turn to connect our framework with existing sub-fields of RL. We examine existing paradigms in these sub-fields from two major perspectives: Scalability and Robustness. Sparse data scenarios require more intelligent ways to use limited experiences. In contrast, abundant data scenarios might suffer from data quality since they might be generated from noisy and often unreliable sources.
@@
measures how methods scale with the increasing problem complexity in terms of the size of the state and action spaces, complex dynamics, noisy reward signals, and longer task horizons. On the one hand, methods might specifically require low dimensional spaces and might not scale so well with increasing the size of these spaces, and on the other, some methods might be overkill for simple problems but better suited for large spaces.
@@
In the following subsections, we cover sub-fields of RL that lie at different areas of the Scalability and Robustness space. Each subsection covers an existing sub-field and its challenges. We then present some examples in which our framework can bolster further research and practice in these fields. Finally, we collate this discussion into takeaways that can be combined into specific areas of further research. These are shown in the blue boxes at the end of each section.
@@ 7.1 Offline RL
Offline Reinforcement Learning (also known as batch RL) (?) involves learning from a fixed dataset without further interaction with the environment. This approach can benefit when active exploration is costly, risky, or infeasible. Consequently, such methods are highly data-dependent due to their reliance on the collected dataset, and they do not generalize well due to the limitations of the pre-collected data. The three dominant paradigms in Offline RL – Behavior Cloning (?), Q-Learning (?), and Sequence Modelling (?) – uniformly degrade in performance as the state-space increases (?). Offline RL has challenges, including effectively overcoming distributional shifts and exploiting the available dataset. Structural decomposition can be crucial in addressing these challenges in different ways, as summarized in the following paragraphs.
@@
Task decomposition allows learning individual policies or value functions for different subtasks, which could potentially leverage the available data more effectively. For example, modular decomposition utilizing portfolios of policies for individual modules using the corresponding subset of the data might be more sample-efficient than learning a single policy for the entire task. Task decompositions, thus, open up new avenues for developing specialized algorithms that effectively learn from limited data about each subtask while balancing the effects of learning different subtasks. Practitioners can leverage such decompositions to maximize the utility of their available datasets by training models that effectively handle specific subtasks, potentially improving the overall system’s performance with the same dataset.
@@
The structural information could potentially help mitigate the effect of distributional shifts. For instance, if some factors are less prone to distributional shifts in a factored decomposition, we could focus more on those factors during learning. This opens up venues for gaining theoretical insights into the complex interplay of structural decompositions, task distributions, and policy performance. On the other hand, practical methods for environments where distributional shifts are standard could leverage structural decomposition to create more robust RL systems.
@@
Structural decomposition can be used to define auxiliary tasks that facilitate learning from the dataset. For instance, in a relational decomposition, we could define auxiliary tasks that involve predicting the relationships between different entities, which could help in learning a valuable representation of the data. Using the proposed framework, researchers can explore how to define meaningful auxiliary tasks that help the agent learn a better representation of the environment. This could lead to new methods that efficiently exploit the available data by learning about these auxiliary tasks. Practitioners can design auxiliary tasks based on the specific decompositions of their problem. For example, if the task has a clear relational structure, auxiliary tasks that predict the relations between different entities can potentially improve the agent’s understanding of the environment and its overall performance.
Research Area 1: Structured Offline RL Use a Factored or Relational decomposition to create abstractions that can help with distribution shift and auxiliary interpretability. Implement a Modular design with each module targeting a specific sub-problem, improving Scalability. Employ policy reuse by policy portfolios learned for sub-problems across tasks. If sufficient interaction data is available, employ data augmentation strategies for counterfactual scenarios using latent models.